[The following information applies to the questions displayed below.]

Cane Company manufactures two products called Alpha and Beta that sell for $185 and $150, respectively. Each product uses only one type of raw material that costs $8 per pound. The company has the capacity to annually produce 119,000 units of each product. Its average cost per unit for each product at this level of activity are given below:

Alpha Beta
Direct materials $ 40 $ 24
Direct labor 33 28
Variable manufacturing overhead 20 18
Traceable fixed manufacturing overhead 28 31
Variable selling expenses 25 21
Common fixed expenses 28 23
Total cost per unit $ 174 $ 145

The company considers its traceable fixed manufacturing overhead to be avoidable, whereas its common fixed expenses are unavoidable and have been allocated to products based on sales dollars.

11. How many pounds of raw material are needed to make one unit of each of the two products? (Alpha/Beta)

12. What contribution margin per pound of raw material is earned by each of the two products? (Round your answers to 2 decimal places.) (Alpha/Beta)

13. Assume that Cane’s customers would buy a maximum of 93,000 units of Alpha and 73,000 units of Beta. Also assume that the raw material available for production is limited to 227,000 pounds. How many units of each product should Cane produce to maximize its profits? (Alpha/Beta)

14. Assume that Cane’s customers would buy a maximum of 93,000 units of Alpha and 73,000 units of Beta. Also assume that the raw material available for production is limited to 227,000 pounds. What total contribution margin will it earn?

15. Assume that Cane’s customers would buy a maximum of 93,000 units of Alpha and 73,000 units of Beta. Also assume that the raw material available for production is limited to 227,000 pounds. If Cane uses its 227,000 pounds of raw materials, up to how much should it be willing to pay per pound for additional raw materials? (Round your answer to 2 decimal places.)
Answer #1

11) Pound needed for one unit

Alpha beta
Pound needed for per unit 40/8 = 5 24/8 = 3

12) Contribution margin per pound

Alpha Beta
Selling price 185 150
Direct material 40 24
Direct labor 33 28
Variable manufacturing overhead 20 18
Variable selling expenses 25 21
Contribution margin per unit 67 59
Pound per unit 5 3
Contribution margin per pound 13.4 19.67

13) Optimum mix

Hour Unit
Beta 73000*3 = 219000 73000
Alpha 8000 8000/5 = 1600
Total 227000

14) Maximum contribution margin = (73000*59+1600*67) = $4414200

15) Highest price = 13.40+8 = 21.40